OpenStudy (anonymous):

Determine the number of real solutions of x2 + 5x + 8 = 0. Show all your steps. (Hint: Use the discriminant.)

OpenStudy (anonymous):

a = 2, b = 5, c = 8, put that into the discriminant we get x = -5 + 5^2 -4(2)(8) and since it = 0 it has 1 real number and 1 rational root.

OpenStudy (anonymous):

5^2 -4(2)(8) = 25 - 64 = -39

OpenStudy (anonymous):

I have no idea what I am doing....

OpenStudy (anonymous):